The one-letter sequence is: SHIRK

a) What is the isoelectric point?

b) What is the average charge on the population of peptide macromolecules at pH = 6.0?

c) What is the average charge on the population of peptide macromolecules at pH = 9.2?

d) What is the average charge on the population of peptide macromolecules at pH = 10.5?
Answer #1

a) Isoelectric point is the pH at which a particular molecule carries no net electrical charge or is electrically neutral in statistical mean.

b) Qpeptide, Qp= (+1)/(1+10(pH-8.5)) + (+2)/(1+10(pH-10.5)) + (-1)/(1+10-(pH-3.2)) + (-4)/(1+10-(pH-3.9)) ------(A)

use value of pH= 6.0 in above equation number A:

Therefore, Qp= -1.97 (at pH=6.0)

c) Qp at pH= 9.2

Use value of pH= 9.2 in equation A:   

Qp= -2.93 (at pH= 9.2)

d) Qp at pH= 10.5

Use value of pH= 10.5 in equation A:

Qp= -2.99 (at pH= 10.5)